# Quantum resolution limit of long-baseline imaging using distributed entanglement

### v1

@@ Abstract
It has been shown that shared entanglement between two telescope sites can in principle be used to localize a point source by mimicking the standard phase-scanning interferometer, but without physically bringing the light from the distant telescopes together. In this paper, we show that a receiver that employs spatial-mode sorting at each telescope site, combined with pre-shared entanglement and local quantum operations can be used to mimic the most general multimode interferometer acting on light collected from the telescopes. As an example application to a quantitative passive-imaging problem, we show that the quantum-limited precision of estimating the angular separation between two stars can be attained by an instantiation of the aforesaid entanglement based receiver. We discuss how this entanglement assisted strategy can be used to achieve the quantum-limited precision of any complex quantitative imaging task involving any number of telescopes. We provide a blueprint of this general receiver that involves quantum transduction of starlight into quantum memory banks and spatial mode sorters deployed at each telescope site, and measurements that include optical detection as well as qubit gates and measurements on the quantum memories. We discuss the relative contributions of local mode sorting at telescope sites vis-a-vis distributed entanglement-assisted interferometry, to the overall quantum-limited information about the scene, based on the ratio of the baseline distance to the individual telescope diameter.
@@ I Introduction
In recent years, novel techniques inspired by quantum information theory have emerged in the field of super-resolution imaging, a field of study aimed at increasing the resolution of an imaging system by overcoming the Rayleigh criterion present in diffraction-limited systems, described in Lord Rayleigh (1879). One very promising technique in this regard is spatial-mode demultiplexing (SPADE), which involves spatial mode sorting of the light entering the imaging system. The light is thus measured in a nontrivial spatial mode basis instead of direct detection of its raw intensity profile on a standard imaging screen, which amounts to measuring the incoming photons in the position (or a ‘pixel mode’) basis. This approach has received considerable attention in recent years following Tsang et al.’s seminal paper on the problem of estimating the separation between two equally bright incoherent point sources Tsang et al. (2016a). Using the classical and quantum Cramér-Rao bounds, they show that while this problem suffers from Rayleigh’s curse when direct detection is employed, a measurement in the image-plane Hermite-Gauss (HG) basis when we have a Gaussian point spread function (PSF), can bypass this limit and attain the optimal estimation precision allowed by the laws of physics which is given by the so-called quantum Fisher Information (QFI). This result was generalized to arbitrary aperture shapes by Kerviche et al. Kerviche et al. (2017) and Rehacek et al. R̆ehác̆ek et al. who independently showed that an optimal SPADE basis for the two equally bright point separation problem can be generated by taking derivatives of increasing orders of the point spread function (PSF) and then performing a Gram-Schmidt orthogonalization process. This yields the sinc-bessel modes for a hard aperture in one dimension. More theory on this has been developed in Tsang et al. (2016a); Ang et al. (2017); Tsang (2017, 2019a); Yu and Prasad (2018), such as extending this to two and three dimensions. Other works have considered estimation of multiple parameters in a scene Prasad and Yu (2019); Řehaček et al. (2017); R̆ehác̆ek et al. (2018); Prasad (2020a); Bisketzi et al. (2019), size estimation of extended objects Dutton et al. (2019); Prasad (2020b), moment estimation of arbitrary scenes Tsang (2019a, b); Zhou and Jiang (2019), object discrimination Lu et al. (2018); Huang and Lupo (2021); Grace and Guha (2022a), adaptively estimating the locations and brightnesses of several point sources of light Bao et al. (2021); Matlin and Zipp (2022); Lee et al. (2022a), finding the precise location of an extended object or an array of point sources Sajjad et al. (2021), or estimating different parameters in two stages such as first finding the centroid (i.e., average position) and then the separation between two points Tsang et al. (2016a); Grace et al. (2020).
@@
These developments have also inspired some works on the application of quantum information theory techniques to the use of multiple optical receivers or apertures to perform long-baseline interferometry Lupo et al. (2020); Wang et al. (2021); Bojer et al. (2022); Sajjad et al. (2024). Such a collection of telescopes roughly emulates a single giant telescope with a size equal to the largest distance (baseline) between two nodes of the array, and its working principle consists of interferometrically exploiting the path-length difference of light arriving at two different observation points Campbell (1987); Law (2000); Monnier (2003). In such a receiver, information can be extracted in two ways in principle. There is the information contained in the path difference between the light collected at different sites, and then there is also spatial resolution residing in the multi-spatial-mode light collected by each telescope. While most of the works in this field from a quantum information theory point of view have focused on the former alone and mostly taken individual receivers to be small, both effects have been considered in full detail in Sajjad et al. (2024), as the first such treatment. They discuss three broad sets of approaches for combining the light collected by different telescopes. First, when the telescopes are close to, e.g., within a few meters of, one another, we can use at each site a carved out piece of a giant parabola spanning the entire baseline, and focus the light on to a common imaging screen where it can be measured. For larger distances where this is not feasible, we can use a spatial mode sorter at each site, couple the individual local first-few orders of information-rich orthogonal spatial modes into single mode fibers (SMFs), bring them to a central location, process them in a linear interferometer, and measure the outputs with shot-noise-limited photon detectors. This is mathematically equivalent to carrying out a spatial mode measurement on a (hypothetical) giant imaging screen spanning the entire baseline Sajjad et al. (2024). A related mathematically-equivalent possibility is to employ multi-spatial-mode ‘light pipes’ or vacuum beam guides Huang et al. (2023) that carry the entire multi-mode light from each telescope to a central location, feed them into a linear interferometer made of bulk elements (e.g., beam splitters) that preserve the multi-spatial-mode information, and measure the outputs of the interferometer after spatial mode sorting at each output. One could use the light-pipe based interferometric receiver as above, but use bucket detectors at each output of the interferometer rather than spatial-mode-resolved detection. Such a strategy would reap the benefit of the baseline, but not extract scene information contained in local high-order spatial modes.
@@
Both approaches, i.e., using either single mode fibers or light pipes, become challenging to realize for baselines beyond a few hundred meters. Losses in single-mode fiber (for the fiber approach) and accurately transporting the multimode light without distortion and loss (for the light pipe approach) become prohibitive, not to mention the physical impracticality of laying light pipes over large geographical distances through oceans, mountains and other forms of complicated terrain. For larger distances, to attain the fundamental resolution precision afforded by a long baseline system, the alternative is to rely on the (future) quantum internet—built using quantum repeaters and satellite-assisted links—that delivers error-corrected high-fidelity entanglement at high rates among the telescope sites, combined with local spatial-mode-resolved light collection and local interferometric measurements at each telescope site.
@@
The simplest example that brings to light the aforesaid issues of quantum limit of a long baseline system, while quantifying the information contributions from mode sorting at each site vis-a-vis entanglement based interferometry, is to consider a two telescope set up for the problem of estimating the angular separation between two stars. At each telescope location, we collect $K$ orthogonal spatial modes with the modes indexed $q=0,\ldots,K-1$ and couple each into single mode fibers (SMFs), e.g., using a free-space-to-fiber coupled multi-plane light conversion (MPLC) mode sorter. Then, for each mode index $q$, we bring the respective SMFs from the two sites to a central location and interfere them in a 50-50 beam splitter, followed by detecting both outputs (of each of the $K$ beamsplitters), where the outputs of the $q$-th beam splitter are the sum and difference of the $q$-th spatial mode signals from the two telescopes. It has been shown in Ref. Sajjad et al. (2024) that in the large $K$ limit, this pairwise interferometric scheme approaches the quantum Fisher information (QFI) limit of precision for estimating the separation between two uniformly bright point sources in one dimension.
@@
In this paper, we describe how such a pairwise measurement scheme can be implemented using shared entanglement in place of single mode fibers and beam splitters by leveraging and building upon: (1) the quantum protocols for long baseline interferometry proposed by Gottesman et al. (2012), (2) its extension to employ solid-state spin-based quantum memories proposed by Khabiboulline et al. (2019a), and (3) modal imaging techniques for super-resolution with a single telescope Tsang et al. (2016b). We also show how our approach can be generalized to carrying out a measurement on any arbitrary set of linear combinations of the different modes collected by any number of telescopes. This way, in principle, we can build the most general entanglement-based receiver and tackle any parameter estimation in a quantitative imaging problem Lee et al. (2022a); Grace and Guha (2022b); Sajjad et al. (2024), for a long-baseline system with any number of telescopes.
@@
In the most general case, let us say we have $n$ telescopes indexed $\alpha=0,\dots,n-1$ where modes $q=0,\ldots,K-1$ are collected at each one of them (since it is impossible to collect an infinite number of modes). And we bring these modes to a central location through single mode fibers where they are combined in a linear interferometer with a desired unitary transformation, which will have $nK$ input modes and $nK$ output modes. Such a linear optical interferometer with any arbitrary unitary transformation can be created using a total of $nK(nK-1)/2$ Mach-Zehnder interferometers (MZIs) Clements et al. (2016). Since a Mach-Zehnder interferometer comprises two 50-50 beam splitters and two phase shifters, a total of $nK(nK-1)$ 50-50 beam splitters and the same number of (tunable) phase shifters suffice to realize a general linear optical unitary. This means that if we can create the equivalence of the action of a 50-50 beam splitter (acting on a single-photon state spread across the two modes) by using shared entanglement, in principle we can carry out the same protocol $nK(nK-1)$ times along with the same number of phase transformations to reproduce the effect of an arbitrary beam splitter circuit with shared entanglement, but without co-locating light from the $n$ telescopes. But, that is exactly what Gottesman et al.’s proposal Gottesman et al. (2012) shows how to do.
@@
Therefore, local spatial-mode sorting at each telescope site, locally-applied single-mode phases and entanglement assisted nonlocal realizations of two-mode 50-50 beamsplitters (see Fig. 1) can be used to compile the most general nonlocal multimode linear-optical pre-detection interferometry, thus helping attain quantum-limited resolution for quantitative imaging problems discussed above, but limited only by the geometry of, and the light collected by the $n$-telescope baseline.
@@
Before closing this section, we should mention how this paper fits into the overall picture in terms of existing proposals on the use of shared entanglement for baseline astronomy Gottesman et al. (2012); Tang et al. (2016); Khabiboulline et al. (2019a, b); Chen et al. (2023); Czupryniak et al. (2023). In the pioneering protocol for entanglement-based interferometry proposed by Gottesman et al., they consider the problem of estimating the phase difference between two telescopes sites for an incoming photon. Say we have $n=2$ sites where we re-label $A$ and $B$ for $\alpha=0$ and $\alpha=1$, respectively. A photon arriving from a star will be in the state $\left(|1_{A},0_{B}\rangle+e^{i\theta}|0_{A},1_{B}\rangle\right)/\sqrt{2}$, where $|1_{A},0_{B}\rangle$ is the state for the photon arriving at telescope $A$, and $|0_{A},1_{B}\rangle$ if it arrives at location $B$, with $\theta$ being a phase difference arising from the path difference for the photon between the two telescopes. This phase will naturally depend on the position of the star and the distance between the two telescopes, and therefore, in concrete terms, estimating $\theta$ amounts to determining the unknown location of a star emitting monochromatic light. In traditional baseline interferometry, we can estimate $\theta$ by bringing the light to a common location through ‘light pipes’, put a tunable phase delay in one of these two signals, feed the two lightpipes into a 50-50 beam splitter, and measure the output intensities. As the tunable phase delay matches $\theta$, one of the two outputs becomes zero with totally destructive interference, and the other output contains all the light. Gottesman et al.’s proposal mimics this technique by using pre-shared Bell states in the optical ‘single-rail’ basis, i.e., $(|0,1\rangle+|1,0\rangle)/\sqrt{2}$, which can be generated by splitting a single photon in a 50-50 beamsplitter, distributed between the telescopes sites reliably via quantum repeaters, local interferometric measurements at each telescope site, and a tunable phase at one telescope site. The main limitation of this scheme is that this entanglement must be provided at an unfeasible rate with the technology available today, and that known repeater architectures—including all-photonic ones Pant et al. (2017)—do not deliver photonic Bell states in the single-rail basis suitable for the aforesaid direct local interferometric measurements. Alternative approaches Khabiboulline et al. (2019a, b); Czupryniak et al. (2022) to phase estimation utilize atomic spin memories in cavities assisted with a binary code to logarithmically compress the required number of pre-shared Bell states. This results in exponentially decreased resource requirements compared to the previous proposal in Gottesman et al. (2012), as well as natural compatibility with atomic-qubit-based quantum repeater architectures Guha et al. (2015); Dhara et al. (2021, 2023); Lee et al. (2022b).
@@
While these contributions have nicely laid down the ground work for entanglement based interferometry, phase estimation between two sites only allows us to estimate the position of a single object. Moreover, a recent suite of work on quantitative passive imaging shows that single-source localization is not a good proxy for imaging more complex scene features, and that most photon efficient schemes must involve multimode, potentially adaptive, interferometry prior to detection Lee et al. (2022a); Grace and Guha (2022b). In order to carry out more complex quantitative imaging and/or parameter estimation tasks, we need to be able to create an entanglement based equivalent of the conventional multiple telescope interferometric receiver where all the light received at different sites is brought through single mode fibers or ‘light pipes’ to a central location, and fed into a linear interferometer whose outputs are then measured for the number of photon clicks. In Ref. Padilla et al. (2024), we studied different alternatives of performing this and to create the entanglement-based equivalent of any multiple telescope receiver with spatial mode sorting.
@@
The paper is organized as follows. In Section II, we explain our physical set up with two telescopes and two uniformly bright point sources. In Section III, we write down the joint quantum state of the incoming light. We then dive into our protocol in Section IV, where the first step is to encode the state of the incoming photons at each telescope site on to memory qubits with the above-mentioned logarithmic compression. In Section V, we explain the decoding step where we carry out the measurements to determine the state of an incoming photon in terms of the combinations of the spatial modes collected at both sites. In Section VI, we discuss how this approach attains the quantum limit of performance for the uniformly bright two-point separation problem. Section VII discusses how our protocol generalizes to any arbitrary measurement with any number of apertures. In Section VIII, we conclude the paper with discussions pertaining to future work.
II Our physical set-up
We consider the following set-up: we want to resolve two thermal light emitters, and assume them to be monochromatic and of equal brightness. Our array consists of two telescopes labeled as systems $A$ and $B$, respectively. For convenience, let us label these systems with $\alpha\in\{A,B\}$. Each site has an aperture that we assume to be symmetrically positioned in the aperture plane. In the case of imaging sources emitting light in optical frequencies (from UV to IR), the incoming light can be modeled as a weak field. Under this assumption, we can model the incoming light as single-rail (absence or presence of a photon) qubit states among the telescope sites, i.e., at a given site, we either have one photon or none. Now, it is convenient to set a measuring time $T$ and subdivide it into $M$ equal intervals of duration $t=T/M$. We can take the probability of detecting a single photon in our two telescope system in each temporal mode (i.e. interval) to be $\epsilon$. The state of a single photon coming in will be a superposition of the states where it arrives at either of the telescopes. One implication of the scheme from Khabiboulline et al. (2019a) being adopted here is that it is only able to encode a single photon during the integration time with an approximate number of time bins $M\approx 1/\epsilon$. It is then desirable to apply such protocol $N$ times to collect $N$ photons in the ideal scenario. By employing a SPADE at each receiver site and pointing it to the average position known as the centroid, we can take the sources to be located at ±$\pm\theta$ in the object plane (see Fig. 2). For this, we require prior knowledge of the centroid, which can be achieved reasonably accurately by employing direct detection Tsang et al. (2016a); Sajjad et al. (2021); Grace et al. (2020). Our goal is to estimate $\theta$ as accurately as possible within the quantum limit Helstrom (1976). This can be achieved by measuring the one-photon signal to be in the symmetric or anti-symmetric combinations of each spatial mode at the two telescope sites Sajjad et al. (2024) by employing shared entanglement. More explicitly, such a set-up would thus measure the light entering the two-aperture system in the basis

### v2

@@ Abstract
It has been shown that shared entanglement between two telescope sites can in principle be used to localize a point source by mimicking the standard phase-scanning interferometer, but without physically bringing the light from the distant telescopes together. In this paper, we show that a receiver that employs spatial-mode sorting at each telescope site, combined with pre-shared entanglement and local quantum operations can be used to mimic the most general multimode interferometer acting on light collected from the telescopes. As an example application to a quantitative passive-imaging problem, we show that the quantum-limited precision of estimating the angular separation between two stars can be attained by an instantiation of the aforesaid entanglement based receiver. We discuss how this entanglement assisted strategy can be used to achieve the quantum-limited precision of any complex quantitative imaging task involving any number of telescopes. We provide a blueprint of this general receiver that involves quantum transduction of starlight into quantum memory banks and spatial mode sorters deployed at each telescope site, and measurements that include optical detection as well as qubit gates and measurements on the quantum memories. We discuss the relative Fisher-information contributions of local mode sorting at telescope sites vis-a-vis distributed entanglement-assisted interferometry, to the overall quantum-limited information about the scene, based on the ratio of the baseline distance to the individual telescope diameter.
@@ I Introduction
In recent years, novel techniques inspired by quantum information theory have emerged in the field of super-resolution imaging, a field of study aimed at increasing the resolution of an imaging system by overcoming the Rayleigh criterion present in diffraction-limited systems, described in Lord Rayleigh (1879). One very promising technique in this regard is spatial-mode demultiplexing (SPADE), which involves spatial mode sorting of the light entering the imaging system prior to detection. The information-bearing light is thus measured in a nontrivial spatial mode basis instead of direct detection of its raw intensity profile on a standard imaging screen that amounts to measuring the incoming photons in the position (or a ‘pixel mode’) basis. This approach has received considerable attention in recent years following Tsang et al.’s seminal paper on the problem of estimating the separation between two equally bright incoherent point sources Tsang et al. (2016a), at the ultimate precision allowed for by quantum estimation theory. Using the classical and quantum Cramér-Rao bounds, they show that while this problem suffers from Rayleigh’s curse when direct detection is employed, a measurement in the image-plane Hermite-Gauss (HG) basis when we have a Gaussian point spread function (PSF), can bypass this limit and attain the optimal estimation precision allowed by the laws of physics, which is given by the so-called quantum Fisher Information (QFI). This result was generalized to arbitrary aperture shapes by Kerviche et al. Kerviche et al. (2017) and Rehacek et al. R̆ehác̆ek et al. who independently showed that an optimal SPADE basis for the two equally bright point separation problem can be generated by taking derivatives of increasing orders of the point spread function (PSF) and then performing a Gram-Schmidt orthogonalization process—henceforth termed the PSF-adapted (PAD) basis spatial mode basis. The PAD basis for a hard-aperture pupil in one dimension comprise of the Sinc-Bessel modes, for example. More theory on this has been developed in Tsang et al. (2016a); Ang et al. (2017); Tsang (2017, 2019a); Yu and Prasad (2018), such as extending this to two and three dimensions. Other works have considered estimation of multiple parameters in a scene Prasad and Yu (2019); Řehaček et al. (2017); R̆ehác̆ek et al. (2018); Prasad (2020a); Bisketzi et al. (2019), size estimation of extended objects Dutton et al. (2019); Prasad (2020b), moment estimation of arbitrary scenes Tsang (2019a, b); Zhou and Jiang (2019), object discrimination Lu et al. (2018); Huang and Lupo (2021); Grace and Guha (2022a), adaptively estimating the locations and brightnesses of several point sources of light Bao et al. (2021); Matlin and Zipp (2022); Lee et al. (2022a), finding the precise location of an extended object or an array of point sources Sajjad et al. (2021), or estimating different parameters in two stages such as first finding the centroid (i.e., average position) and then the separation between two points Tsang et al. (2016a); Grace et al. (2020).
@@
These developments have also inspired some works on the application of quantum information theory techniques to the use of multiple optical receivers or apertures to perform long-baseline interferometry Lupo et al. (2020); Wang et al. (2021); Bojer et al. (2022); Sajjad et al. (2024). Such a collection of telescopes roughly emulates a single giant telescope with a size equal to the largest distance (baseline) between two nodes of the array, and its working principle consists of interferometrically exploiting the path-length difference of light arriving at two different observation points Campbell (1987); Law (2000); Monnier (2003). In such a receiver, information can be extracted in two ways in principle. There is the information contained in the path difference between the light collected at different sites, and then there is also spatial resolution residing in the multi-spatial-mode light collected by each telescope. While most of the works in this field from a quantum information theory point of view have focused on the former alone and mostly taken individual receivers to be small, both effects have been considered in full detail in Sajjad et al. (2024), as the first such treatment. They discuss three broad sets of approaches for combining the light collected by different telescopes. First, when the telescopes are close to, e.g., within a few meters of, one another, we can use at each site a carved out piece of a giant parabola spanning the entire baseline, and focus the light on to a common imaging screen where it can be measured, e.g., by a single SPADE whose modal basis is matched to the problem at hand and the PSF of the effective compound aperture system. For larger distances where this is not feasible, we can use a spatial mode sorter at each site, couple the individual local first-few orders of information-rich orthogonal spatial modes into single mode fibers (SMFs), bring them to a central location, process them in a linear interferometer, and measure the outputs with shot-noise-limited photon detectors. This is mathematically equivalent to carrying out a SPADE measurement on a (hypothetical) giant image plane spanning the entire baseline Sajjad et al. (2024) discussed above. A related equivalent possibility is to employ multi-spatial-mode ‘light pipes’ or vacuum beam guides Huang et al. (2023) that carry the entire multi-mode light from each telescope to a central location, feed them into a linear interferometer made of bulk elements (e.g., beam splitters) that preserve the multi-spatial-mode information, and measure the outputs of the interferometer after spatial mode sorting at each output. One could use the light-pipe based interferometric receiver as above, but use bucket detectors at each output of the interferometer rather than spatial-mode-resolved detection. Such a strategy would reap the benefit of the baseline, but not extract scene information contained in local high-order spatial modes.
@@
Both approaches to baseline interferometry, i.e., using either single mode fibers or light pipes, become challenging to realize for baselines beyond a few hundred meters. Losses in single-mode fibers (for the fiber approach) and accurately transporting the multimode light without distortion and loss (for the light pipe approach) become prohibitive, not to mention the physical impracticality of laying light pipes over large geographical distances through oceans, mountains and other forms of complicated terrain. For larger distances, to attain the fundamental resolution precision afforded by a long baseline system, the alternative is to rely on the (future) quantum internet—built using quantum repeaters and satellite-assisted links—that delivers error-corrected high-fidelity entanglement at high rates among quantum-memory banks deployed at the telescope sites, combined with local spatial-mode-resolved light collection and local interferometric measurements at each telescope site.
@@
The simplest example that brings to light the aforesaid issues of quantum limit of a long baseline system, while quantifying the information contributions from mode sorting at each site vis-a-vis entanglement based interferometry, is to consider a two telescope set up for the problem of estimating the angular separation between two stars. At each telescope location, we collect $K$ orthogonal spatial modes with the modes indexed $q=0,\ldots,K-1$ and couple each into single mode fibers (SMFs), e.g., using a free-space-to-fiber coupled multi-plane light conversion (MPLC) mode sorter. Then, for each mode index $q$, we bring the respective SMFs from the two sites to a central location and interfere them in a 50-50 beam splitter, followed by detecting both outputs (of each of the $K$ beamsplitters), where the outputs of the $q$-th beam splitter are the sum and difference of the $q$-th spatial mode signals from the two telescopes. It has been shown in Ref. Sajjad et al. (2024) that in the large $K$ limit, this pairwise interferometric scheme approaches the quantum Fisher information (QFI) limit of the precision for estimating the separation between two uniformly bright point sources in one dimension using two telescopes.
@@
In this paper, we first describe how such a pairwise interferometric receiver can be implemented non-locally using pre-shared entanglement in place of single mode fibers and beam splitters. We do this by leveraging and building upon: (1) the quantum protocols for long baseline interferometry proposed by Gottesman et al. (2012), (2) its extension to employ solid-state spin-based quantum memories proposed by Khabiboulline et al. (2019a), and (3) modal imaging techniques for super-resolution with a single telescope Tsang et al. (2016b). We then show how our approach can be generalized for carrying out parameter estimation in any complex quantitative imaging problem Lee et al. (2022a); Grace and Guha (2022b); Sajjad et al. (2024) at the quantum precision limit corresponding to a long-baseline system with any number of telescopes, using pre-shared multi-site entanglement and local operations at each site. We do this by mimicking a photon measurement in any arbitrary linear combination of optical modes collected at the telescope sites by a series of pairwise two-mode beamsplitter operations realized nonlocally via bi-partite pre-shared entangled Bell pairs.
@@
In the most general case, let us say we have $n$ telescopes indexed $\alpha=0,\dots,n-1$ where local spatial modes indexed $q=0,\ldots,K-1$ are collected and sorted at each one of them. We bring these $K$ modes, coupled into single mode optical fibers, to a central location where they are combined in a linear interferometer in a desired unitary transformation, i.e., one with $nK$ input modes and $nK$ output modes. Such a linear optical interferometer realizing an arbitrary unitary mode transformation can be created using a total of $nK(nK-1)/2$ Mach-Zehnder interferometers (MZIs) Clements et al. (2016). Since a Mach-Zehnder interferometer comprises two 50-50 beam splitters and two single-mode phase shifters, a total of $nK(nK-1)$ 50-50 beam splitters and the same number of (tunable) single-mode phase shifters suffice to realize a general linear optical unitary. This means that if we can create the equivalence of the action of a 50-50 beam splitter (acting on a single-photon state spread across the two nonlocal modes on which the beamsplitter acts) by using pre-shared entanglement, in principle we can carry out $nK(nK-1)$ such nonlocal beamsplitters along with the same number of (local) single mode phases to reproduce the effect of an arbitrary beam splitter circuit consuming up to $nK(nK-1)$ pre-shared Bell pairs between quantum memories held at pairs of telescope sites, but without ever co-locating light from the $n$ telescopes. In this paper, we show how to realize the building block of the above argument—a two-input nonlocal beamsplitter acting on a photon spread across the two nonlocal modes—by first loading the photon mode at each site into a quantum memory Khabiboulline et al. (2019a) along with local quantum gates acting on those memories, and a preshared Bell pair among the two sites, via a sequence of local quantum gates.
@@
Therefore, local spatial-mode sorting at each telescope site, locally-applied single-mode phases and entanglement-assisted nonlocal realizations of two-mode 50-50 beamsplitters (see Fig. 1) can be used to compile the most general nonlocal multimode linear-optical pre-detection interferometry, thus helping attain quantum-limited resolution for quantitative imaging problems discussed above, limited only by the geometry of, and the light collected by the $n$-telescope baseline.
@@
Before closing this section, we should mention how this paper fits into the overall picture in terms of existing proposals on the use of shared entanglement for baseline astronomy Gottesman et al. (2012); Tang et al. (2016); Khabiboulline et al. (2019a, b); Chen et al. (2023); Czupryniak et al. (2023). In the pioneering protocol for entanglement-based interferometry proposed by Gottesman et al., they consider the problem of estimating the phase difference between a pair of optical modes collected at the two telescopes sites carrying one incoming photon. Say we have $n=2$ sites where we re-label telescope sites $A$ and $B$ for $\alpha=0$ and $\alpha=1$, respectively. A photon arriving from a star is in the state $\left(|1_{A},0_{B}\rangle+e^{i\theta}|0_{A},1_{B}\rangle\right)/\sqrt{2}$, where $|1_{A},0_{B}\rangle$ is the state for the photon arriving at telescope $A$, and $|0_{A},1_{B}\rangle$ if it arrives at location $B$, with $\theta$ being a phase difference arising from the path difference for the photon between the two telescopes. This phase will naturally depend on the position of the star and the distance between the two telescopes, and therefore, in concrete terms, estimating $\theta$ amounts to determining the unknown location of a star emitting monochromatic light. In traditional baseline interferometry, we can estimate $\theta$ by bringing the light to a common location through ‘light pipes’, put a tunable phase delay in one of these two signals, feed the two lightpipes into a 50-50 beam splitter, and measure the output intensities. As the tunable phase delay matches the equivalent optical phase shift of $\theta$, one of the two output intensities becomes zero due to fully destructive interference, and the other output contains all the light. Gottesman et al.’s proposal mimics this technique by using pre-shared Bell states in the optical ‘single-rail’ basis, i.e., $(|0,1\rangle+|1,0\rangle)/\sqrt{2}$, where $|1\rangle$ denotes a single-photon Fock state of a mode and $|0\rangle$ is vacuum, which can be generated by splitting a single photon in a 50-50 beamsplitter, distributed between the telescopes sites (reliably, say via quantum repeaters), local interferometric measurements at each telescope site, and a tunable phase at one telescope site. The main limitation of this scheme is that this entanglement must be provided at an unfeasible rate (at the modes-per-second corresponding to the optical bandwidth of the ‘science photons’ being collected) with the technology available today, and that known repeater architectures—including all-photonic ones Pant et al. (2017)—do not deliver photonic Bell states in the single-rail basis suitable for the aforesaid direct local interferometric measurements. They would deliver Bell states in the ‘dual-rail’ photonic-qubit basis, converting which to single rail is not easy. Alternative approaches were proposed Khabiboulline et al. (2019a, b); Czupryniak et al. (2022) for the aforesaid phase estimation utilizing atomic spin based quantum memories in cavities, assisted with a binary-code logarithm compression of the quantum state of one science photon spread across $M$ temporal modes into $\log_{2}(M)$ qubit memories, thereby resulting in an exponential reduction in the pre-shared Bell-state resource requirements compared to the previous proposal in Gottesman et al. (2012), as well as lending itself to a natural compatibility with atomic-qubit-based quantum repeater architectures Guha et al. (2015); Dhara et al. (2021, 2023); Lee et al. (2022b).
@@
While these contributions have nicely laid down the ground work for entanglement based interferometry, estimation of the phase difference between an incoming science photon between two telescope sites allows us only to estimate the position of a single point object. One could argue that with repeated use of this protocol, one could image any general scene. However, a suite of recent work on quantitative passive imaging problems shows that single-point-source localization is not a good proxy for imaging more complex scene features, and that the most photon efficient schemes—ones whose performance attains respective quantum limits of those quantitative imaging problems—must involve multimode, potentially adaptive, interferometry prior to detection Lee et al. (2022a); Grace and Guha (2022b). In order to carry out more complex quantitative imaging and/or parameter estimation tasks, we need to be able to create an entanglement based equivalent of a multiple-telescope interferometric receiver where all the light received at different sites is brought through single mode fibers or ‘light pipes’ to a central location, and fed into a linear interferometer whose outputs are then measured for the number of photon clicks Sajjad et al. (2024).
@@
The paper is organized as follows. In Section II, we explain our problem set up with two telescopes and two uniformly bright point sources, and summarize the main results of the paper. In Section III, we write down the joint quantum state of the incoming information-bearing light. We then dive into our nonlocal entanglement-assisted measurement protocol in Section IV, where the first step is to encode the state of the incoming photons at each telescope site on to memory qubits with the above-mentioned logarithmic compression idea. In Section V, we explain the decoding step where we carry out the measurements to determine the state of an incoming photon in terms of the combinations of the spatial modes collected at both sites. In Section VI, we discuss how this approach attains the quantum Fisher-information (QFI) limit of estimation precision for the uniformly bright two-point separation problem. Section VII discusses how our protocol generalizes to any arbitrary measurement with any number of telescopes. In Section VIII, we conclude the paper with discussions pertaining to future work.
II Problem set up and main results
We consider the problem of estimating the half-separation $\theta$ between two equally-bright incoherently-radiating quasi-monochromatic point sources located at angular positions ±$\pm\theta$ (see Fig. 2). We will assume that a sufficiently-accurate prior knowledge of the centroid of the two sources has been obtained by direct detection using a small fraction of the incoming photons during the integration time Tsang et al. (2016a); Sajjad et al. (2021); Grace et al. (2020). Our imaging system consists of a two-telescope array labeled $\alpha\in\{A,B\}$, centered at ±$\pm\beta$, each of diameter $\delta$ (see Fig. 2). The number of temporal modes $M$ of the incoming light contained within an integration time of $T$ (seconds) is: $M\approx WT$, where $W$ (Hz) is the optical bandwidth around the center frequency. We will consider imaging sources emitting light in optical frequencies (from UV to IR), in which regime the incoming light can be modeled as a weak thermal field. We will take the mean photon number per temporal mode $\epsilon$, in the incoming light collected by the two telescopes together, to be much less than one. We consider an integration time window $T$ commensurate with $M\approx 1/\epsilon$, such that there is roughly one photon but no more that arrives within that time window across both telescopes. We will adopt the compressive cavity-mediated light-matter interaction scheme from Ref. Khabiboulline et al. (2019a) to load the quantum state of $1$ photon spanning $M$ temporal modes, into $\log_{2}(M)$-qubit quantum memory registers located at each telescope site. The problem we consider is to estimate $\theta$ as accurately as possible—per the quantum-estimation-theory limit Helstrom (1976)—given a total of photons $N$ collected across both telescope sites.
Ref. Sajjad et al. (2024) considered this problem, and showed that the Quantum Fisher Information (QFI) of $\theta$ (the inverse of which gives a tight bound on the minimum-permissible variance in estimation of $\theta$ when $N$ is large) is given by:
${\mathcal{K}}=\frac{4\pi^{2}N}{3\sigma^{2}}\left(3r^{2}+1\right),$ (1)
where $r=2\beta/\delta$ and $\sigma=2\pi/\delta$ is the Rayleigh separation in wavelength-normalized unity-magnification-ratio units. Ref. Sajjad et al. (2024) also showed that—provided the light collected by both apertures could be made available to be (processed and) detected in a common location with no loss—a receiver structure that achieves the QFI would employ spatial mode sorters at each aperture to sort sufficiently many (say $K\geq 1$) mutually-orthogonal spatial modes in the PAD bases (discussed in previous section) of each aperture, followed by projecting the two-mode state of the light collected—in each temporal mode, across both apertures in the $q$-th spatial mode, for each $q$, $10\leq q\leq K-$ 1—into the symmetric or anti-symmetric mutually-orthogonal two-mode single-rail photonic-qubit Bell states: